# Orbital decay candidates reconsidered: WASP-4 b is not decaying and Kepler-1658 b is not a planet

### v3

@@ 5.1 Radial Velocities
The RV data are shown in the upper left panel of Figure 4. The peak-to-peak variations of about 1,000 m/s have been attributed to the hot Jupiter. Beyond this, as noted by Vissapragada et al. (2022), there is no evidence for a long-term trend. The constraint is weaker than it may appear, though, because the TRES and NEID data can be shifted relative to each another by an arbitrary constant. The primary leverage on any RV trend comes from the consistency between the earliest three TRES velocities and those obtained seven years later.
@@ 5.2 Transit Times
We re-analyzed the TESS light curves from Sectors 41, 54, and 55, and incorporated newly TESS data from Sectors 74, 75, 81, and 82. Because single transits are barely detectable in the TESS photometry, we abandoned attempts to measure individual transit times. Instead, we grouped the data into four multi-transit blocks: Sector 41, Sectors 54 & 55, Sectors 74 & 75, and Sectors 81 & 82. Within each block, the light curve was fitted assuming strict periodicity, with the key parameter being the transit time of the event closest to the block center. There was no requirement for strict periodicity across different blocks. This approach offers greater robustness at the expense of sensitivity to short-term variations. Vissapragada et al. (2022) used a similar procedure to analyze the Kepler data, treating the observing quarters as independent blocks. We adopted without modification their reported transit times from Kepler and from WIRC.
@@ 5.3 Analysis
We scanned $10^{4}$ trial values of $P_{2}$, log-uniformly distributed between $7$ and $10^{5}$ days. For each choice of $P_{2}$, we optimized the other parameters and recorded the value of $\chi^{2}$. The optimal solution was found at $P_{2}\approx 2500$ days, with $\chi^{2}=187$ with 52 degrees of freedom — a poor fit. We proceeded under the tentative assumption that the RV uncertainties had been underestimated, which seemed plausible since precise RV extraction codes are not optimized for broad-lined stars such as Kepler-1658 ($v\sin i\approx 34$ km/s). By adding terms in quadrature of $\sigma_{\rm TRES}=180$ m/s and $\sigma_{\rm NEID}=100$ m/s, the $\chi^{2}$ of the best-fit model was lowered to equal the number of degrees of freedom.
@@ 5.4 Problems with the planet hypothesis
Seeking corroborating evidence, we re-examined the Kepler photometry. A median light curve was constructed from all the 1-min data, phase-folded using the transit period, and phase-averaged to 2.3-min cadence (124 data points) to speed up further analysis. The standard deviation outside of transits was 22 parts per million (ppm), which we adopted as the per-point uncertainty. Fitting the light curve with the Mandel & Agol (2002) transit model gave $\chi^{2}=134.5$ with 117 degrees of freedom. While this is statistically acceptable, there are structured patterns in the residuals, especially during ingress and egress (Figure 6).
@@ 5.5 Hierarchical triple-star model
  • the dilution factor of $\sim$$10^{2}$ inferred from the Kepler light curve,
@@
Consistency with the observed rate is achieved if the eclipsing binary is located $a\sim 50$ AU from the F star. At the Gaia-determined distance of 788 parsecs (Gaia Collaboration et al., 2023) [^13] https://gaia.ari.uni-heidelberg.de/singlesource.html the implied angular separation is $0.06$ arcsec, consistent with being unresolved.
@@ 7 Acknowledgments
We thank Shreyas Vissapragada, Dan Huber, and Özgür Baştürk for feedback on the manuscript. J.N.W. gratefully acknowledges support from the NASA TESS project and a NASA Keck PI award. The work described in this paper was based in part on observations with the NEID instrument on the WIYN 3.5m telescope at Kitt Peak National Observatory. Kitt Peak is a facility of NSF’s NOIRLab, managed by the Association of Universities for Research in Astronomy (AURA). The WIYN telescope is a joint facility of NOIRlab, Indiana University, the University of Wisconsin-Madison, Pennsylvania State University, Purdue University, and Princeton University. NEID was funded by the NASA-NSF Exoplanet Observational Research (NN-EXPLORE) partnership and built by Pennsylvania State University. The NEID archive is operated by the NASA Exoplanet Science Institute at the California Institute of Technology. NN-EXPLORE is managed by the Jet Propulsion Laboratory, California Institute of Technology under contract with the National Aeronautics and Space Administration. The authors thank the NEID Queue Observers and WIYN Observing Associates, as well as the California Planet Search team supporting precise RV acquisition with the Keck I telescope, for their dedicated and skillful execution of the observations.

### v4

@@ 5.1 Radial Velocities
The RV data are shown in the upper left panel of Figure 4. The peak-to-peak variations of about 1,000 m/s have been attributed to the hot Jupiter. Beyond this, as noted by Vissapragada et al. (2022), there is no evidence for a long-term trend. The constraint is weaker than it may appear, though, because the TRES and NEID data can be shifted relative to each other by an arbitrary constant. The primary leverage on any RV trend comes from the consistency between the earliest three TRES velocities and those obtained seven years later.
@@ 5.2 Transit Times
We re-analyzed the TESS light curves from Sectors 41, 54, and 55, and incorporated new TESS data from Sectors 74, 75, 81, and 82. Because single transits are barely detectable in the TESS photometry, we abandoned attempts to measure individual transit times. Instead, we grouped the data into four multi-transit blocks: Sector 41, Sectors 54 & 55, Sectors 74 & 75, and Sectors 81 & 82. Within each block, the light curve was fitted using the same basic method as Ivshina & Winn (2022a), but assuming strict periodicity, with the key parameter being the transit time of the event closest to the block center. There was no requirement for strict periodicity across different blocks. This approach offers greater robustness at the expense of sensitivity to short-term variations. Vissapragada et al. (2022) used a similar procedure to analyze the Kepler data, treating the observing quarters as independent blocks. We adopted without modification their reported transit times from Kepler and from WIRC.
@@ 5.3 Analysis
We scanned $10^{4}$ trial values of $P_{2}$, log-uniformly distributed between $7$ and $10^{5}$ days. For each choice of $P_{2}$, we optimized the other parameters and recorded the value of $\chi^{2}$. The optimal solution was found at $P_{2}\approx 2500$ days, with $\chi^{2}=187$ with 52 degrees of freedom — a poor fit. We proceeded under the tentative assumption that the RV uncertainties had been underestimated, which seemed plausible since precise RV extraction codes are not optimized for broad-lined stars such as Kepler-1658 ($v\sin i\approx 34$ km/s). By adding jitter terms of $\sigma_{\rm TRES}=180$ m/s and $\sigma_{\rm NEID}=100$ m/s in quadrature with the formal uncertainties, the $\chi^{2}$ of the best-fit model was lowered to equal the number of degrees of freedom.
@@ 5.4 Problems with the planet hypothesis
Seeking corroborating evidence, we re-examined the Kepler photometry. A median light curve was constructed from all the 1-min data: the data were phase-folded using the transit period, and phase-averaged to 2.3-min cadence (124 data points) to speed up further analysis. The standard deviation outside of transits was 22 parts per million (ppm), which we adopted as the per-point uncertainty. Fitting the light curve with the Mandel & Agol (2002) transit model gave $\chi^{2}=134.5$ with 117 degrees of freedom. While this is statistically acceptable, there are structured patterns in the residuals, especially during ingress and egress (Figure 6).
@@ 5.5 Hierarchical triple-star model
  • the dilution factor of $\epsilon\sim 10^{-2}$ inferred from the Kepler light curve,
@@
Consistency with the observed rate is achieved if the eclipsing binary is located $a\sim 50$ AU from the F star. At the Gaia-determined distance of 788 parsecs (Gaia Collaboration et al., 2023) [^14] https://gaia.ari.uni-heidelberg.de/singlesource.html, the implied angular separation is $0.06$ arcsec, consistent with being unresolved.
@@ 6 Discussion
Postscript.—After this manuscript was submitted and posted on the arXiv, several colleagues provided further evidence supporting the false-positive interpretation of Kepler-1658 b. S. Vissapragada found that the $J$-band WIRC light curve is consistent with the $\approx$0.3% transit depth predicted in Section 5. S. Mahadevan and C. Cañas pointed out that Fleming et al. (2015) had previously reported a discrepancy between the RV amplitudes measured at optical and near-infrared wavelengths.
@@ 7 Acknowledgments
We thank Shreyas Vissapragada, Dan Huber, Özgür Baştürk, Suvrath Mahadevan, and Caleb Cañas for feedback on the manuscript. J.N.W. gratefully acknowledges support from the NASA TESS project and a NASA Keck PI award. The work described in this paper was based in part on observations with the NEID instrument on the WIYN 3.5 m telescope at Kitt Peak National Observatory. Kitt Peak is a facility of NSF’s NOIRLab, managed by the Association of Universities for Research in Astronomy (AURA). The WIYN telescope is a joint facility of NOIRLab, Indiana University, the University of Wisconsin-Madison, Pennsylvania State University, Purdue University, and Princeton University. NEID was funded by the NASA-NSF Exoplanet Observational Research (NN-EXPLORE) partnership and built by Pennsylvania State University. The NEID archive is operated by the NASA Exoplanet Science Institute at the California Institute of Technology. NN-EXPLORE is managed by the Jet Propulsion Laboratory, California Institute of Technology under contract with the National Aeronautics and Space Administration. The authors thank the NEID Queue Observers and WIYN Observing Associates, as well as the California Planet Search team supporting precise RV acquisition with the Keck I telescope, for their dedicated and skillful execution of the observations.
Some of the data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute and can be accessed via doi:10.17909/19s5-fe67.
Appendix A New transit times and RVs for WASP-4 and WASP-12
Table 2: WASP-4 HIRES relative radial velocities
Time Radial velocity Uncertainty
(BJD) (m/s) (m/s)
2455378.10273 -191.35 2.84
2456152.12045 218.23 3.18
2456168.07375 199.24 3.14
2456207.94924 -60.88 2.83
2456532.09391 193.72 2.69
2458662.10006 -326.8 2.67
2458715.08168 -31.28 2.74
2458723.06334 0.35 3.17
2459189.71567 -1.21 2.77
Table 3: New WASP-4 transit times
Time of conjunction Uncertainty
(BJDTDB) (days)
2460909.86707 0.00034
2460911.20587 0.00034
2460912.54389 0.00032
2460913.88185 0.00033
2460915.22090 0.00033
2460916.55882 0.00032
2460917.89716 0.00036
2460919.23551 0.00035
2460923.24982 0.00034
2460924.58793 0.00035
2460925.92689 0.00036
2460928.60266 0.00032
2460929.94090 0.00035
2460931.27952 0.00034
2460932.61711 0.00037
Note. — Based on TESS Sector 96.
Table 4: WASP-12 NEID relative radial velocities
Time Radial velocity Uncertainty
(BJD) (m/s) (m/s)
2460356.79680 19038.4 7.0
2460364.79782 18806.5 7.2
2460370.79990 19217.2 6.9
2460374.81484 19033.9 9.0
2460381.69721 19229.3 7.4
2460382.74067 19246.1 10.3
2460383.71488 19234.2 8.0
2460388.73542 18793.7 10.1
2460390.77048 18883.9 10.7
2460392.71535 19113.1 12.0
2460415.63353 19142.8 9.8
2460439.63265 19154.9 11.0
2460606.94516 18812.6 6.2
2460607.90961 18934.7 7.3
2460608.80832 19141.8 9.4
2460609.95971 19081.4 6.2
2460691.69033 19231.5 8.5
2460693.79940 19252.7 5.0
2460694.83252 19260.0 5.9
2460706.67271 19119.0 7.2
2460730.63381 19050.3 5.5
2460731.60757 18923.6 5.0
2460732.61346 18834.8 6.1
2460733.83508 18951.2 8.4
2460734.67715 18793.2 9.1
2460737.67836 19042.7 6.3
2460739.67883 19239.4 5.0
2460743.65539 18973.1 5.2